Find the center (h,k) of the circle with the equation x^2+y^2-6y=7.
(h,k)=(_,_)
a. 0,9
b.-3,9
c.0,3
d.9,16
Thanks for your help..........

Answer by kiru_khandelwal(79)

To find centre of the circle with equation x^2+y^2-6y = 7
Convert the equation to the form
(x-h)^2 + (y-k)^2 = r^2
where h and k are coordinates of the centre of the circle and r is the radiys of the circie
so x^2 + Y^2 -6y +9 = 7+9
=> x^2 +(y-3)^2 = 16
=> (x-0)^2 + (y-3)^2 = 4^2
so centre of the circle is (0,3) and radius is 4